### v1

# UniGenX: Unified Generation of Sequence and Structure with Autoregressive Diffusion

Abstract
Unified generation of sequence and structure for scientific data (e.g., materials, molecules, proteins) is a critical task. Existing approaches primarily rely on either autoregressive sequence models or diffusion models, each offering distinct advantages and facing notable limitations. Autoregressive models, such as GPT, Llama, and Phi-4, have demonstrated remarkable success in natural language generation and have been extended to multimodal tasks (e.g., image, video, and audio) using advanced encoders like VQ-VAE to represent complex modalities as discrete sequences. However, their direct application to scientific domains is challenging due to the high precision requirements and the diverse nature of scientific data. On the other hand, diffusion models excel at generating high-dimensional scientific data, such as protein, molecule, and material structures, with remarkable accuracy. Yet, their inability to effectively model sequences limits their potential as general-purpose multimodal foundation models. To address these challenges, we propose UniGenX, a unified framework that combines autoregressive next-token prediction with conditional diffusion models. This integration leverages the strengths of autoregressive models to ease the training of conditional diffusion models, while diffusion-based generative heads enhance the precision of autoregressive predictions. We validate the effectiveness of UniGenX on material and small molecule generation tasks, achieving a significant leap in state-of-the-art performance for material crystal structure prediction and establishing new state-of-the-art results for small molecule structure prediction, de novo design, and conditional generation. Notably, UniGenX demonstrates significant improvements, especially in handling long sequences for complex structures, showcasing its efficacy as a versatile tool for scientific data generation.
Keywords— Multimodality, Diffusion, Autoregressive, Science generation
1 Introduction
The ability to accurately model and generate both sequence and structure information within scientific data is a linchpin for advancing artificial intelligence in scientific discovery. This challenge, though critical, presents a significant hurdle, as the intricate interplay between sequence and structure often dictates the fundamental properties and functionalities of materials, molecules, and proteins. Structural information, therefore, serves as the bedrock for comprehending their behavior and designing transformative applications. For instance, the precise arrangement of atoms within a crystal lattice directly influences a material’s mechanical, thermal, and electronic characteristics [1], while the nuanced three-dimensional conformation of a protein governs its biological function through specific molecular interactions [2]. Similarly, the spatial configuration of small molecules exerts a profound influence on their chemical reactivity, pharmacological efficacy, and potential toxicity [3].
In these scientific domains, where precision is paramount, even subtle deviations in atomic coordinates or bond angles can precipitate substantial alterations in properties, rendering a material unsuitable for its intended application or disrupting a molecule’s interaction with critical biological targets [4]. For example, minor inaccuracies in molecular conformations can lead to erroneous predictions of binding affinities in drug discovery pipelines [5] or impede the precise replication of desired material properties in materials science laboratories [6]. Furthermore, the impact of structural inaccuracies is compounded when modeling larger, more complex systems, such as intricate protein complexes, extended periodic materials, or extensive molecular networks, where the complexity of the system amplifies the criticality of precision [7].
The inherent high-dimensional nature of scientific data further exacerbates the complexity of this modeling task. Unlike the symbolic or approximate representations often employed in natural language processing or image analysis, scientific data frequently relies on precise numerical representations, including three-dimensional atomic coordinates, lattice vectors, and energetic properties [8]. These representations are not only exceptionally sensitive to minor perturbations but also exhibit intricate, multiscale dependencies, spanning from atomic-level interactions to macroscopic properties. Consequently, the successful modeling of such data necessitates a profound understanding of both sequence and structure, coupled with the capacity to generate outputs that are both physically plausible and scientifically meaningful [9].
Adding to the complexity is the inherent diversity of scientific data formats across various domains. Materials are typically represented using chemical formulas and periodic lattice structures [10], molecules are characterized by SMILES strings and three-dimensional conformations [11], and proteins are described by amino acid sequences and spatial folding patterns [12]. Each of these representational paradigms encodes unique, complex relationships between sequence and structure, demanding specialized approaches to accurately capture their intricacies. The absence of a unified modeling framework capable of seamlessly handling this diversity significantly impedes the development of general-purpose artificial intelligence systems for scientific discovery.
Consequently, achieving the dual objectives of high precision and broad generalization in the modeling of sequence and structure remains a substantial bottleneck in the field of AI for science. Overcoming this challenge is crucial for advancing applications in diverse domains such as materials design, drug discovery, and protein engineering, where the ability to accurately model and predict the intricate interplay between sequence and structure can unlock transformative innovations [4, 13].
Autoregressive language models (LLMs), exemplified by models such as GPT [14], Llama [15], phi-4 [16], Gemini [17], and DeepSeek [18], have revolutionized the field of language generation and have been successfully extended to multimodal tasks, including image, video, and audio generation, through the utilization of vector quantized embeddings. These models benefit significantly from scaling laws, where increasing model size and data volume leads to improved performance. Their ability to process diverse data as token sequences, through direct tokenization for text or vector quantized embeddings for other modalities, confers both flexibility and generalizability. Furthermore, their capacity to handle long-context inputs enables the generation of complex outputs for both conditional and unconditional tasks. However, these models encounter significant challenges when applied to scientific domains, particularly in the generation of numerically accurate outputs, which is indispensable for tasks such as three-dimensional structure generation.
Conversely, diffusion models [19, 20, 21, 22] have demonstrated exceptional prowess in high-dimensional numerical generation tasks, achieving remarkable success in generating high-resolution images and videos [23, 24, 25, 26, 27]. These models typically encode high-dimensional data, such as images with dimensions of $512\times 512\times 3$, into lower-dimensional latent representations, such as $128\times 128\times h$ where $h$ is the hidden dimension, using encoders like variational autoencoders (VAEs). These representations are then patch-wisely tokenized, for example, $32\times 32$ with 1024 tokens, and processed by transformer-based models [28] to predict noise or score values. These tokens are subsequently treated as a concatenated high-dimensional vector ($1024\times h$), upon which the diffusion process is conditioned. While this approach enhances efficiency compared to directly diffusing in the original image space, training diffusion models remains computationally intensive and prone to defects [29], especially when not fully trained. Moreover, diffusion models generally struggle with sequence modeling tasks, limiting their applicability in scenarios requiring flexible multimodal capabilities.
The scientific domain, with its intricate numerical representations and diverse data formats, presents unique challenges for generative models. Even minute deviations in atomic coordinates can drastically alter molecular properties or render generated structures meaningless. While LLMs have shown promise in specific scientific applications [30, 31, 32, 33, 34, 35], they still struggle with unified numerical data modeling. Conversely, diffusion models are successful in generating structures [36, 37, 38, 39, 40, 41], but lack flexible multimodal capabilities.
To address these challenges, we introduce UniGenX, a novel framework integrating autoregressive next-token prediction with conditional diffusion models. This leverages autoregressive flexibility and diffusion precision, overcoming numerical accuracy limitations while maintaining multimodal capability.
Our approach offers several key advantages:
  • Improved numerical accuracy: The conditional diffusion model head enhances numerical precision by leveraging input from the autoregressive module, which simplifies the diffusion process by focusing on the next token.
  • Flexible generation: The autoregressive framework enables the generation of sequences with varying lengths, accommodating diverse scientific tasks.
  • Multimodality: By combining diffusion models with autoregressive modeling, UniGenX supports flexible multimodal data generation, bridging sequence and structure information.
  • State-of-the-Art Performance: Preliminary results demonstrate that UniGenX achieves significant improvements over existing methods across multiple scientific benchmarks.
Existing models often jointly diffuse sequence and structure or use sequences as embeddings. These approaches exacerbate training difficulties or fail to model sequences adequately. UniGenX handles long sequences effectively, demonstrating superior performance in crystal structure prediction for large systems.
We evaluate UniGenX on materials and small molecules, spanning tasks like crystal structure prediction, de novo material generation, conformation generation, and conditional molecule generation. UniGenX significantly outperforms FlowMM [39] on MP-20, Carbon-24, and MPTS-52 benchmarks and achieves state-of-the-art results in small molecule tasks. Trained on a unified dataset, UniGenX demonstrates cross-domain generalization. Incorporating pretraining with NatureLM [35] extends to natural language processing, validating our approach’s versatility.
In summary, UniGenX is a unified generative framework for science, integrating sequence and structure generation. By combining autoregressive and diffusion models, it addresses key limitations, enabling accurate and flexible scientific data generation.
2 Background
2.1 Autoregressive Model
Auto-regressive models have become a cornerstone in deep learning for sequential data modeling. These models operate by factorizing the joint probability distribution of a sequence into a product of conditional probabilities, allowing each token or data point to be predicted based on the preceding ones. Formally, given a sequence
$\mathbf{x}={x_{1},x_{2},\ldots,x_{T}},$
the auto-regressive approach models the probability as
$P(\mathbf{x})=\prod_{t=1}^{T}P(x_{t}\mid x_{1},x_{2},\ldots,x_{t-1}),$
where $x_{i}$ is the $i$-th token or element in the sequence. This property makes them highly effective for tasks such as language modeling, where predicting the next word in a sentence requires contextual understanding of the preceding words. Popular examples include models like the GPT series [14] and LLaMA [15, 42, 43], which primarily utilize multi-head attention-based Transformer [28] decoder architectures to capture long-range dependencies in data.
At the heart of these transformer [28] decoder architectures lies the multi-head attention mechanism. Multi-head attention works by projecting an input sequence into multiple subspaces, allowing the model to attend to different parts of the sequence simultaneously. For a given token, the mechanism computes attention weights that quantify how much focus should be given to each other token in the sequence. These weights are then used to aggregate information from the sequence, creating a context-aware representation of the token.
Specifically, for an input sequence represented as a matrix $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$ (where $T$ is the sequence length and $d_{\text{model}}$ is the model’s dimensionality), the input embeddings are projected into $h$ independent subspaces, where $h$ is the number of attention heads. This is achieved by using separate learned weight matrices for each $i$-th head:
$\mathbf{Q}_{i}=\mathbf{X}\mathbf{W_{Q}}^{(i)},\quad\mathbf{K}_{i}=\mathbf{X}\mathbf{W_{K}}^{(i)},\quad\mathbf{V}_{i}=\mathbf{X}\mathbf{W_{V}}^{(i)},\quad i=1,2,\ldots,h.$
Here, $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$, $\mathbf{V}_{i}$ are called query, key and value respectively. $\mathbf{W_{Q}}^{(i)}$, $\mathbf{W_{K}}^{(i)}$, and $\mathbf{W_{V}}^{(i)}\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$ are the learned weight matrices for the $i$-th attention head, and $d_{k}$ is typically set to $d_{\text{model}}/h$ so that the total computational cost remains constant.
The attention weights are calculated as scaled dot-products between the queries and keys, followed by a softmax operation to normalize them:
$\text{Attention}(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i})=\text{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}_{i},$
where Attention $i\in\mathbb{R}^{T\times d_{k}}$ is the output of the $i$-th attention head, and the division by $\sqrt{d_{k}}$ serves as a scaling factor to prevent the dot product values from becoming too large, which could otherwise lead to small gradients during backpropagation.
The final output is calculated by concatenating the outputs for $h$ heads along the feature dimension and followed by a linear transformation:
$\mathbf{O}=\text{Concat}(\text{Attention}_{1},\text{Attention}_{2},\ldots,\text{Attention}_{h})\cdot\mathbf{W_{O}},$
where $\text{Concat}(\cdot)$ denotes the concatenation operation, $\mathbf{W_{O}}\in\mathbb{R}^{(h\cdot d_{k})\times d_{\text{model}}}$. The resulting matrix $\mathbf{O}\in\mathbb{R}^{T\times d_{\text{model}}}$ represents the final output of the multi-head attention mechanism, which is then passed to subsequent layers in the transformer architecture.
The transformer decoder architecture applies multi-head attention in an autoregressive manner by masking future tokens during training. This is achieved by using a causal mask, which prevents the model from attending to tokens that come after the current token in the sequence. This masking ensures that predictions are conditioned only on previously observed tokens, preserving the autoregressive property.
2.2 Diffusion Models
2.2.1 Diffusion Process
A diffusion model is defined by a diffusion process $\{x_{t}\}_{t=0}^{1}$, starts with $x\sim p_{0}$, and evolves to $x_{1}\sim p_{1}$, where $p_{1}$ has a tractable form for efficient sample generation. The dataset of independent samples from the data distribution is available. Let $p(x_{t})$ denote the probability density of $x_{t}$, and $p\left(x_{t}\mid x\right)$ represent the transition kernel from $x$ to $x_{t}$, with $0<t\leq 1$. The diffusion process is modeled as the solution to a stochastic differential equation (SDE) [44]:
$\mathrm{d}x_{t}=f(x_{t},t)\mathrm{d}t+g(t)\mathrm{d}\bm{B}_{t},$ (1)
where $\bm{B}_{s}$ is a standard Wiener process, $f(\cdot,t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift coefficient of $x(t)$, and $g(\cdot):\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient of $x(t)$.
By reversing the diffusion process, starting from samples $x_{1}\sim p_{1}$, where $p_{1}$ is usually a Gaussian distribution, it is possible to recover samples $x\sim p_{0}$. The reverse of a diffusion process is also a diffusion process, governed by the reverse-time SDE [22]:
$\mathrm{d}x_{t}=\left[f(x_{t},t)-g(t)^{2}\nabla_{x_{t}}\log p_{t}(x_{t})\right]\mathrm{d}t+g(t)\mathrm{d}\overline{\bm{B}},$ (2)
where $\overline{\bm{B}}$ is another standard Wiener process, i.e., Brownian motion from $t=1$ to $t=0$.
As the diffusion process is a stochastic process with the Brownian motion stochastic term, the SDE (1) of the forward process implies the following relationship
[…]
where the conditional probability $p(x_{t}|x)=C\exp(-\frac{\|x_{t}-\alpha_{t}x\|^{2}}{\sigma_{t}^{2}})$, and $C$ is the normalization constant satisfying $\int_{x_{t}}C\exp{-\frac{\|x_{t}-\alpha_{t}x\|^{2}}{\sigma_{t}^{2}}}dx_{t}=1$. Equivalently, we have
$x_{t}=\alpha_{t}x+\sigma_{t}\epsilon.$ (4)
Here $\alpha_{t}$ and $\sigma_{t}$ are linear and variance variable, depending on the time $t$; $\epsilon$ is a random variable. In the diffusion model, $\alpha_{t}$ and $\sigma_{t}$ are given by designed schemes for VE and VP cases [22]; the random variable $\epsilon$ is a standard Gaussian noise variable, i.e., $\epsilon\sim N(0,I)$. In our work, we use the DDPM setting for training and the VP scheme for sampling [21, 44]. We also test the diffusion choice of EDM [45]. We obtained similar performance leveraging DDPM and EDM respectively, and listed the results as an ablation study in Appendix E.3.2.
2.2.2 Diffusion Targets
Different from conventional supervised learning, the training targets in the training loss of the diffusion model are not obvious most of the time. With the same noising process, the training targets differ between different models, such as $\epsilon$ model [21], denoising model [46], and score model [20].
Here we would show that all of these targets have connected forms. Recall that adding noise by $x_{t}=\alpha_{t}x+\sigma_{t}\epsilon$, i.e.,
$p(x_{t}|x)=Ce^{-\frac{-\|x_{t}-\alpha_{t}x\|^{2}}{2\sigma_{t}^{2}}}.$ (5)
By the definition of the score function $s(x_{t})\equiv\nabla_{x_{t}}\log p(x_{t})$ and through a direct calculation, we have
[…]
where we finally obtain
$s(x_{t})=\mathbb{E}_{x}[\nabla_{x_{t}}\log p(x_{t}|x)|x_{t}]=\mathbb{E}_{x}[-\frac{x_{t}-\alpha_{t}x}{\sigma_{t}^{2}}|x_{t}]=-\frac{x_{t}-\alpha_{t}\mathbb{E}[x|x_{t}]}{\sigma_{t}^{2}}.$ (7)
By Theorem 1, the score of a distribution, $\nabla_{x_{t}}\log p_{t}(x_{t})$, can be estimated by training a score-based model $s_{\bm{\theta}}$ using score matching [20] with the training target $\nabla_{x_{t}}\log p(x_{t}|x)$. To approximate $\nabla_{x_{t}}\log p(x_{t})$, a time-dependent score-based model $s_{\theta}(x_{t},t)$ can be trained using a continuous generalization of the score matching objective [22]:
[…]
where $\lambda:[0,1]\rightarrow\mathbb{R}^{+}$ is a weighting function, $U[0,1]$ is a uniform distribution in $[0,1]$, $x\sim p_{0}$ and $p_{0}$ is the data distribution.
From Eq. (7), one can also approximate the training target with a different model, such as the denoising model $x_{\theta}$ with the target $x$ [46], and the $\epsilon$ model $\epsilon_{\theta}$ [21] with the target $\frac{x_{t}-\alpha_{t}x}{\sigma_{t}}=\epsilon$, with a similar training loss as the score model. Hence, the optimal solution of the models (denoted as $\theta^{*}$) approaches:
$s_{\theta^{*}}\to\mathbb{E}_{x}[\nabla_{x_{t}}\log p(x_{t}|x)|x_{t}];\quad\epsilon_{\theta^{*}}\to\mathbb{E}_{x}[\frac{x_{t}-\alpha_{t}x}{\sigma_{t}}|x_{t}];\quad x_{\theta^{*}}\to\mathbb{E}_{x}[x|x_{t}].$ (8)
These models connect to each other in the optimal results given by Eq. (7), and have the following forms
$s_{\theta^{*}}(x_{t})$ $=-\frac{x_{t}-\alpha_{t}x_{\theta^{*}}(x_{t})}{\sigma_{t}^{2}}$ (9)
$s_{\theta^{*}}(x_{t})$ $=-\frac{\epsilon_{\theta^{*}}(x_{t})}{\sigma_{t}}.$
Therefore, when training diffusion models using an $\epsilon$ model or denoising model, the sampling process can be aligned with score-based models via Eq. (9) and utilize the reverse stochastic differential equation (2). In this work, we adopt the $\epsilon$ model proposed in DDPM [21] for training and employ the VP scheme [22] for sampling.
2.3 Related Work
Recent research has explored the integration of autoregressive models with diffusion-based generative techniques to overcome the limitations of each approach and enhance overall generation capabilities. By leveraging diffusion models as a complementary mechanism within autoregressive frameworks, researchers aim to improve the precision and efficiency of tasks like continuous embedding generation, which traditionally posed challenges for autoregressive methods due to their reliance on vector quantization. One prominent line of research involves adding diffusion models as lightweight components on top of existing autoregressive language models. These models enhance precision in the generation of continuous representations by combining the strengths of both approaches. For example, the Masked Autoregressive (MAR) model [47] and LatentLLM [48] incorporate diffusion-based approaches to refine hidden representations and improve the generation of continuous embeddings. The NOVA model [49] exemplifies the success of this hybrid approach by integrating diffusion methods into an autoregressive framework for multimodal tasks. The result is a system that can handle complex generative tasks with higher precision and versatility, showcasing the potential of combining autoregressive and diffusion techniques to push the boundaries of generative modeling. These works are most closely related to ours, but require advanced encoders for other modalities (such as images) to generate hidden embeddings, followed by alignment with language in a latent space.
3 Methods
UniGenX leverages the strengths of both autoregressive (AR) and diffusion models, achieving a natural unification. The diffusion head addresses the numerical precision limitations of general AR models by operating in a continuous vector space for numerical data. Conversely, the AR next-token prediction provides effective conditioning for the diffusion head, which diffuses only a low-dimensional variable (e.g., three dimensions for molecular 3D structures), significantly simplifying the diffusion training process. Our design seamlessly integrates symbolic (words) and numerical (numbers) data across scientific domains without sacrificing scientific precision. This section details the key features of the UniGenX architecture.
3.1 Sequentialize All for Multi-Domain/Task Compatibility
Figure 1: Sequentialization approach combines symbolic (words) and numerical data (numbers), separated by special tokens that identify the category of each data element. As illustrated in the figure for small molecules and materials, SMILES strings or chemical formulas are treated as words and delimited by special tokens such as $<$bos$>$ (beginning of SMILES) or $<$bof$>$ (beginning of formula). Corresponding coordinates are treated as numbers, delimited by tokens such as $<$boc$>$ (beginning of coordinates). This method provides a general framework applicable to diverse scientific data formats.
The success of multimodal GPT-like models in integrating language and images demonstrates the potential of sequence alignment for addressing diverse problems across various domains. To achieve alignment with language—a naturally sequential modality—images are encoded into a latent space using a pretrained encoder and subsequently aligned with language embeddings within this shared latent space. This process enables general comprehension and facilitates generation in both language and image modalities. This approach highlights the flexibility of sequentialization across diverse domains and tasks, a principle validated by previous research and applications of GPT models.
However, scientific data presents distinct challenges in sequentialization compared to data from general domains like language, images, and videos. Scientific data exhibits significantly greater complexity and diverse standards across different disciplines. For instance, small molecules are frequently represented using SMILES strings, encoding atoms and chemical bonds [50], while periodic materials are often described by chemical formulas [10], accompanied by atomic coordinates. Proteins, DNA, and data types like energies and forces also utilize specialized, compact representations. This inherent domain specificity and the deep integration of expert knowledge within these representations make establishing a unified standard exceptionally challenging. Consequently, developing a universal representation capable of aligning structures, energies, forces, and other numerical properties with language-like atomic or molecular descriptions is exceedingly difficult. Furthermore, scientific data inherently contains a large amount of numerical data that is highly sensitive to numerical precision. The aforementioned challenges preclude the direct transfer of sequentialization methods from general domains to scientific applications.
To fully leverage the flexibility of multimodality, we propose a simple yet effective sequentialization approach: representing all data—formulas, coordinates, energies, forces, and so on—as a single sequence segmented by special tokens. This sequentialization enables framing all generation tasks as next-token prediction problems, thereby harnessing the scalability and flexibility inherent in autoregressive models for scientific applications. And we will show that this approach can handle different scientific domains with different tasks.
Figure 1 depicts our sequentialization scheme, which unifies symbolic (words) and numerical data (numbers) into a single sequence delineated by special tokens. For small molecules, the $<$bos$>$ token marks the onset of the SMILES string, while $<$boc$>$ signals the commencement of its coordinate sequence. A parallel approach is applied to materials, with $<$bof$>$ initiating the chemical formula and $<$boc$>$ the coordinate sequence. The $<$eos$>$ token is appended to signify the sequence’s termination. This methodology facilitates the encoding of diverse scientific data into a unified sequence format and is adaptable to other data types, including protein and DNA sequences.
This sequentialization method offers a simple and effective way to represent diverse scientific data across various domains and tasks. However, traditional sequence models struggle to handle numerical data (numbers) with the same efficacy as symbolic data (words), particularly when numerical precision is crucial. To address this challenge, our model incorporates a diffusion strategy combined with a sequence model that effectively handles both numerical and symbolic data within a sequence modeling framework. This approach will be detailed in the following subsection.
3.2 Bridge the Gap: Unified Sequence Modeling of Words and Numbers
Figure 2: The scheme of the UniGenX framework for a unified generation. We present a novel autoregressive diffusion framework for science generation, surpassing traditional LLMs and structural prediction models. By combining next-token prediction and conditional diffusion, our approach offers enhanced flexibility, overcomes the difficulties of training of the traditional jointly diffusion, and the numerical accuracy issues that have limited the application of LLMs in high-precision scientific fields.
[…]
This design treats words and numbers equally within the autoregressive framework but employs distinct heads to model the discrete and continuous spaces, respectively. The majority of parameters reside within the autoregressive module, responsible for next-token prediction, while the diffusion head is significantly lighter. This is because the autoregressive module captures most of the contextual information in the predicted token, leaving only a low-dimensional variable for the diffusion process. This approach differs substantially from traditional joint diffusion training and consequently requires far fewer parameters and time steps for training the diffusion head. The diffusion head operates on a single variable at a time, for example, the three-dimensional coordinates or forces of a specific atom. Consequently, the diffusion training is agnostic to the semantic meaning of the variable; whether it represents coordinates, energy, or force is irrelevant to the diffusion process itself. Therefore, a single diffusion head can be used for training all numerical data without loss of precision. Therefore, our model exhibits significant flexibility across diverse domains and tasks that can be represented as sequences of words and numbers.
Our model does not incorporate explicit inductive biases for equivariance or invariance. This design choice grants significant flexibility in handling diverse tasks and scientific domains with varying symmetry requirements. We demonstrate through our results that learning specific equivariant or invariant properties can be effectively achieved through data augmentation, thus preserving the scalability of the underlying transformer architecture.
4 Experiments
To evaluate UniGenX, we conduct experiments across various tasks in the materials and small molecule domains. In this work, we present two model configurations of different sizes: UniGenX(100M) and UniGenX(400M), where the transformer decoder backbones contain 100M and 400M parameters, respectively. See Table S5 for detailed model configurations.
4.1 Material
For material domain, we evaluate UniGenX on crystal structure prediction and de novo generation task.
4.1.1 Crystal Structure Prediction
In this task, we performed finetuning and evaluations on three mainstream benchmarks. The MP-20 dataset is a subset of the Materials Project [51]. It encompasses all materials with a unit cell atomic count less than 20, and most properties can achieve DFT accuracy. Carbon-24 [52] contains 10k carbon materials, which share the same composition, but have different structures. Similar to the MP-20 dataset, it covers materials composed only of carbon atoms with a unit cell atomic count less than 24. The MPTS-52 dataset includes time-based splits for cross-validation and benchmarking generative models, allowing up to 52 atoms in the structures.
And the results will be compared with three baselines which are all previous state-of-the-art methods. CDVAE [53] is a VAE-based framework designed for generating pure crystals. It is modified in [38] to perform the CSP task by replacing its original normal prior for generation with a parametric prior conditioned on the encoding of the specified composition. DiffCSP [38] is based on diffusion model. It jointly generates lattice vectors and fractional coordinates to learn the structure distribution of stable crystals. The latest SOTA metheod is FlowMM [39], which uses Riemannian Flow Matching to create distributions over fractional atomic coordinates, unit cells, and atomic types, ensuring invariance to translations, rotations, and permutations inherent in crystals. While our model does not incorporate explicit inductive biases for equivariance or invariance, we demonstrate through our results that these properties can be effectively learned through data augmentation, preserving the scalability of the underlying transformer architecture and providing flexibility across diverse tasks and scientific domains with varying symmetry requirements.
[…]
4.1.2 De Novo Generation
As a unified model, UniGenX inherently possesses the capability of de novo generation. Following [39], we let the model finetuned on MP-20 generate 10,000 materials. Different from CSP task, It start with only begin token <bos> and generating the compositions and structures through sampling and diffusion.
We assess the generation performance using three metrics: Validity, Coverage, and Property Statistics. Validity measures the correctness of the predicted crystals. Coverage evaluates the similarity between the test set and the generated samples. Property Statistics calculates properties such as density and the number of elements.
4.2 Molecule
For small molecule domain, we evaluate UniGenX on conformation generation and conditional generation tasks.
4.2.1 Conformation Generation
Adopting the data splits defined by DMCG [36], we conducted training and evaluation on the large-scale GEOM-QM9 and GEOM-Drugs datasets. The GEOM dataset [55] comprises three-dimensional conformations for over 37 million molecules, each annotated with experimental data and high-precision Density Functional Theory (DFT) energy values. GEOM-QM9 and GEOM-Drugs are subsets of this comprehensive dataset. Notably, GEOM-QM9 offers a richer conformational diversity compared to the traditional QM9 dataset [56], incorporating intermediate states in addition to ground states, which is advantageous for downstream tasks such as molecular and protein-molecule docking.
Among the comparative methods, GEOMOL [57] employs message passing neural networks (MPNNs) and SE(3)-invariance to predict local atomic 3D structures and torsion angles, facilitating deterministic assembly of complete conformers. ConfGF [58] is a gradient-based model that optimizes molecular 3D structures by learning interatomic distance gradient fields and utilizing score matching for training. DMCG [36], the current state-of-the-art, directly predicts atomic 3D coordinates through a rotation, translation, and permutation-invariant loss function and an iterative refinement architecture.
For evaluation, given molecule $x$ with $N_{x}$ conformations in the test set, we generate $2N_{x}$ conformations, following [58]. Let $\mathbb{S}_{g}$ and $\mathbb{S}_{r}$ represent the sets of generated and ground truth conformations, respectively. Conformation deviation is quantified using GetBestRMS from the RDKit package, with root-mean-square deviation denoted as $\mathrm{RMSD}(R,\hat{R})$. The recall-based coverage (COV) and matching (MAT) scores are defined as follows:
$\mathrm{COV}(\mathbb{S}_{g},\mathbb{S}_{r})=\frac{1}{|\mathbb{S}_{r}|}\begin{vmatrix}\begin{Bmatrix}R\in\mathbb{S}_{r}|\mathrm{RMSD}(R,\hat{R})<\delta,\exists\hat{R}\in\mathbb{S}_{g}\end{Bmatrix}\end{vmatrix}$ (10)
$\mathrm{MAT}(\mathbb{S}_{g},\mathbb{S}_{r})=\frac{1}{|\mathbb{S}_{r}|}\sum_{R\in\mathbb{S}_{r}}{min}_{\hat{R}\in\mathbb{S}_{g}}\mathrm{RMSD}(R,\hat{R})$ (11)
[…]
$\mathrm{COV}\text{-}\mathrm{P}(\mathbb{S}_{g},\mathbb{S}_{r})=\frac{1}{|\mathbb{S}_{g}|}\begin{vmatrix}\begin{Bmatrix}{\hat{R}\in\mathbb{S}_{g}}|\mathrm{RMSD}(R,\hat{R})<\delta,\exists R\in\mathbb{S}_{r}\end{Bmatrix}\end{vmatrix}$ (12)
$\mathrm{MAT}\text{-}\mathrm{P}(\mathbb{S}_{g},\mathbb{S}_{r})=\frac{1}{|\mathbb{S}_{g}|}\sum_{{\hat{R}\in\mathbb{S}_{g}}}{min}_{R\in\mathbb{S}_{r}}\mathrm{RMSD}(R,\hat{R})$ (13)
4.2.2 Property Prediction
We extended our evaluation to include molecular property prediction alongside conformation generation, a task that involves predicting properties from generated conformations [55]. Following [36], we randomly selected 30 molecular SMILES from the test set of the GEOM-QM9 dataset. For each molecule, we generated 50 distinct conformations. Using the Psi4 quantum chemistry software package [60], we calculated energy levels, HOMO, and LUMO characteristics for both generated and reference conformations. We then derived ensemble averages for energy ($\overline{E}$), minimum energy ($E_{\text{min}}$), average HOMO-LUMO gap ($\Delta\overline{\epsilon}$), and the gap extremities ($\Delta\epsilon_{\text{min}}$ and $\Delta\epsilon_{\text{max}}$) based on the conformational attributes of each molecule. Mean absolute error (MAE) was used to quantify property discrepancies between generated and reference conformations. Given our model’s superior performance in predicting these statistical quantities, depicted in Sec. 5.2, such as average energy and average HOMO-LUMO gap, we further compared the distribution of sampled results from UniGenX with the ground-truth conformations. To our knowledge, this is the first metric of distribution comparison performed for this task.
4.2.3 Conditional Generation
We also explored our model’s conditional generation capabilities. Instead of employing guidance in the diffusion model, UniGenX utilizes a straightforward approach: prepending the training data sequence with the condition property and its value, marked by a special token like $<$bobulk$>$ for the bulk property. This enables simple prompting at inference, where conditional generation is achieved by providing the property-value pair and the sequence generation special token. Data processing examples are detailed in Appendix C.5. In this section, using conditional molecule generation on the QM9 benchmark as an example, we demonstrate UniGenX’s strong performance and note its easy extensibility to different domains.
Following [61], the QM9 dataset is divided into two halves, the first half is used to train the classifier, while the second half is used to train our model. For the classifier, we use the training weights provided by [62] to achieve a fair comparison. Unlike other diffusion models, we wrap the condition into the sequence. Only one special token is needed for wrapping. Refer to Appendix for more details.
The properties include $\alpha(Bohr^{3})$, the ability of a molecule to become polarized under an external electric field; energy gap $\Delta\varepsilon(meV)$, the energy difference between HOMO and LUMO; HOMO $\varepsilon_{H}(meV)$, the energy level of the highest occupied molecular orbital; LUMO $\varepsilon_{L}(meV)$, the lowest unoccupied molecular orbital energy; dipole moment $\mu(D)$, the separation of positive and negative charges within the molecule.
For evaluation, we employed two different sampling methods. The first, introduced in [61], models the distribution of conditions in the training set based on 15. It first samples the number of atoms, $N$ , and then samples the property values corresponding to $N$. The second method, proposed by [62], addresses a potential inductive bias in the original method from [61], where an implicit correlation between molecule size and properties may arise during evaluation. To mitigate this, they propose uniformly sampling property values directly between the minimum and maximum values, as described in 14.
We evaluated our model’s performance aligning both sampling methods. Notably, unlike prior approaches that trained separate models for each property and relied on extensive preprocessing, our model is trained as a unified, all-in-one model, directly utilizing raw data without requiring pre-generation of the number of atoms, N.
$p(\mathcal{M}|\mathbf{x}_{\text{cond}})\propto\sum_{N}p(\mathcal{M}_{N}|N,\mathbf{x}_{\text{cond}})p(N)$ (14)
$p(\mathcal{M}|\mathbf{x}_{\text{cond}})\propto\sum_{N}p(\mathcal{M}_{N}|N,\mathbf{x}_{\text{cond}})p(N)p(\mathbf{x}_{\text{cond}}|N)$ (15)
4.3 Unification training on Material and Molecule
We demonstrated our model’s ability to perform unified training and generation across diverse scientific domains. Specifically, we trained a single model on both molecule and material datasets. As detailed in the methods, UniGenX not only accommodates pure sequences, such as crystal formulas in materials or SMILES strings in small molecules, but also effectively handles numerical tokens, such as atomic coordinates, through its diffusion heads. This showcases UniGenX’s potential for unifying various scientific domains by training a single model across different tasks and datasets.
To differentiate domains, we introduced special tokens for materials and molecules as indicators during training. We trained a 100M parameter model on a combined dataset of 5M material data and 1M QM9 molecule data, denoted as UniGenX(100M). We then fine-tuned this model on material datasets including MP-20, Carbon-24, and MPTS-52, as well as the 1M GEOM-QM9 molecule dataset.
4.4 Combing with Natural Language Model
Building upon our approach, which leverages an autoregressive sequence model backbone and joint training on symbolic and numerical data, UniGenX naturally lends itself to language capabilities. We explored this by loading a pretrained language model and performing instruction tuning on the MP-20 crystal structure prediction task using language prompts. The MP-20 dataset was reconstructed into an instruction tuning dataset, with the data format detailed in Appendix C.6.
We then fine-tuned the pretrained model, which had approximately 1B parameters from NLM [35], on this dataset, demonstrating its superior performance.
5 Results
We evaluated UniGenX on crystal structure prediction and de novo generation in the material domain, conformation generation and conditional generation in the molecular domain, unification and language capacity. Table S2 summarizes the computational resources used, including model size, GPU type, and total GPU days.
5.1 Material Results
MP-20 Carbon-24 MPTS-52
Methods MR(%$\uparrow$ RMSD $\downarrow$ MR(%$\uparrow$ RMSD $\downarrow$ MR(%$\uparrow$ RMSD $\downarrow$
CDVAE[53] 33.90 0.1045 17.09 0.2969 5.34 0.2106
DiffCSP[38] 51.49 0.063 17.54 0.2759 12.19 0.1786
FlowMM[39] 61.39 0.057 23.47 0.4122 17.54 0.1726
NatureLM[35] 61.78 0.044    –    – 30.20 0.084
UniGenX(400M) $\mathbf{67.01}$ $\mathbf{0.037}$ $\mathbf{30.05}$ $\mathbf{0.2286}$ $\mathbf{38.65}$ $\mathbf{0.0657}$
Table 1: Crystal structure prediction results on material benchmarks. “-” indicates unreported results. Our model significantly outperforms the state-of-the-art, with increased performance on longer sequences.
Methods InSteps VS (%$\uparrow$ VC (%$\uparrow$ CR (%$\uparrow$ CP(%$\uparrow$ PW($\rho$) $\downarrow$ PW(Nel) $\downarrow$
CDVAE[53] 5000 $\mathbf{100}$ 86.7 99.15 99.49 0.688 0.278
DiffCSP[38] 1000 $\mathbf{100}$ 83.25 $\mathbf{99.71}$ 99.76 0.35 0.125
FlowMM[39] 250 96.58 83.47 99.48 99.65 0.261 0.107
FlowMM[39] 1000 96.85 83.19 99.49 99.58 0.239 0.083
UniGenX(100M) $\mathbf{200}$ 99.08 $\mathbf{90.12}$ 99.27 $\mathbf{99.95}$ $\mathbf{0.065}$ $\mathbf{0.04}$
Table 2: De novo generation tasks benchmark. IS, VS, VC, CR, CP and PW represent Integration steps, Validity Structural, Validity Composition, Coverage Recall, Coverage Precision and Property Wdist.
Table 1 and Table 2 present the evaluation metrics for baseline methods and our model in Crystal Structure Prediction and de novo generation, respectively. As shown in Table 1, our model significantly outperforms the previous diffusion/flow matching state-of-the-art, FlowMM, achieving match rate increases of 10%, 28%, and 120% on MP-20, Carbon-24, and MPTS-52, respectively. In terms of RMSD, we observe improvements of 35%, 45%, and 62%. Compared to NatureLM [35], a recently published autoregressive model that surpassed FlowMM on the MP-20 and MPTS-52 benchmarks, our model also demonstrates superior performance, establishing a new state-of-the-art across MP-20, Carbon-24, and MPTS-52.
The significant improvement on MPTS-52 is particularly noteworthy. We speculate that this is due to our model’s inherent long-range perception capabilities, a feature often lacking in graph-based representations. The results in Table 1, compared with both diffusion/flow matching and autoregressive models, underscore the importance and effectiveness of combining diffusion and autoregressive models in our approach.
Table 2 shows that UniGenX can generate materials attributed to more reasonable, more precise, and more realistic under the condition of fewer steps. Specifically, on the Property Wdist metric, which is an index on which previous models did not perform well, UniGenX can achieve scores of 0.065 for $\rho$ (density) and 0.04 for Nel (number of electrons) in just 200 steps. Compared to the previously most advanced model, this represents an improvement of 73% and 52%, respectively.
Figure 3: Examples of generated material structures from the crystal structure prediction task. The generated structures closely resemble the ground truth structures; therefore, the latter are omitted for clarity and conciseness.
Figure 4: Examples of generated material structure in de novo generation task
Some examples of crystal structure prediction are shown in Figure 3, and Figure 4 displays some examples of de novo generation.
5.2 Molecular Results
Dataset Large-scale QM9 Large-scale Drugs
[…]
Mean Median Mean Median Mean Median Mean Median
ConfGF[58] 46.23 44.87 0.5171 0.5133 28.23 20.71 1.6317 1.6155
GeoMol[57] 78.28 81.03 0.3790 0.3861 41.46 36.79 1.5120 1.5107
DMCG[36] 90.86 95.36 $\mathbf{0.2305}$ 0.2258 74.57 81.80 $\mathbf{0.9940}$ 0.9454
UniGenX(400M) $\mathbf{91.40}$ $\mathbf{100.00}$ 0.2516 $\mathbf{0.1070}$ $\mathbf{76.94}$ $\mathbf{87.50}$ 1.3394 $\mathbf{0.8631}$
Table 3: Conformation generation: precision-based coverage and matching score. Our model surpasses the state-of-the-art on most metrics.
Dataset Large-scale QM9 Large-scale Drugs
[…]
Mean Median Mean Median Mean Median Mean Median
ConfGF[58] 89.21 95.12 0.2809 0.2837 70.92 85.71 1.0940 1.0917
GeoMol[57] 91.05 95.55 0.2970 0.2993 69.74 83.56 1.1110 1.0864
DMCG[36] $\mathbf{98.34}$ $\mathbf{100}$ 0.1486 0.1340 $\mathbf{96.22}$ $\mathbf{100.00}$ 0.6967 0.6552
UniGenX(400M) 92.67 $\mathbf{100}$ $\mathbf{0.1466}$ $\mathbf{0.0856}$ 92.09 $\mathbf{100.00}$ $\mathbf{0.6536}$ $\mathbf{0.6120}$
Table 4: Conformation generation: Recall-based coverage and matching score
As shown in Table 3 and Table 4, the conformations generated by our model are able to achieve good Coverage and Match score whether based on recall or precision. Overall, the results are comparable to DMCG. It is noticeable that for every metric’s median we outperform the baselines; the remaining ones are only a little lower than DMCG [36].
Methods $\overline{E}$ $E_{\text{min}}$ $\overline{\Delta\epsilon}$ $\Delta\epsilon_{\text{min}}$ $\Delta\epsilon_{\text{max}}$
RDKit 0.8875 0.6530 0.3484 0.5570 0.2399
ConfGF[58] 2.8349 0.2012 0.6903 4.9221 0.1820
GeoMol[57] 4.5700 0.5096 0.5616 3.5083 0.2650
DMCG[36] 0.4324 0.1364 0.2057 1.3229 0.1509
UniGenX(400M) 0.1464 0.0746 0.1004 0.2241 0.1360
Table 5: Mean absolute error of predicted ensemble properties (Unit: eV)
Figure 5: Comparison of energy (A, C) and gap (B, D) property distributions between ground truth and generated conformations. A and C, and B and D, represent alternative visualizations of the same distributions.
The results of property prediction are shown in Table 5. Our method significantly outperforms RDKit, GeoMol, ConfGF and DMCG, which shows the effectiveness of our method. Comparison between the distribution of energies and gaps of ground truth and generated conformations is in Figure 5.
Model $\alpha(Bohr^{3})$ $\epsilon_{H}(mHa)$ $\epsilon_{L}(mHa)$ $\Delta\epsilon(mHa)$ $\mu(D)$
Random 41.00 103.30 121.83 193.36 8.40
EDM[61] 20.15 158.70 166.20 287.00 7.01
LDM-3DG[62] 15.56 54.62 63.08 107.14 6.33
LDM-3DG-GSSL[62] 16.43 55.03 66.53 113.15 9.22
DiGress[63] 9.23 31.98 105.06 90.57 1.49
SeaDAG[64] 8.85 30.91 103.03 89.70 1.33
UniGenX(100M) 2.46 18.40 21.47 55.77 0.99
Table 6: Conditional generation on five quantum properties(unit) evaluation following LDM setting in [62]. Numbers represent the MAE between conditional and oracle-predicted properties [65].
Property $\alpha$ $\Delta\varepsilon$ $\varepsilon_{\mathrm{HOMO}}$ $\varepsilon_{\mathrm{LUMO}}$ $\mu$ $C_{v}$
Units Bohr3 meV meV meV D $\frac{\text{cal}}{\text{mol}}$
QM9 0.10 64 39 36 0.043 0.040
Random 9.01 1470 645 1457 1.616 6.857
EDM[61] 2.76 655 356 584 1.111 1.101
GeoLDM[66] 2.37 587 340 522 1.108 1.025
GeoBFN[67] 2.34 577 328 516 0.998 0.949
Geo2Seq with Mamba[68] 0.46 98 57 71 0.164 0.275
Geo2Seq with GPT[68] 0.53 102 48 53 0.097 0.325
MOL-STRUCTOK[69] 0.33 89 64 62 0.285 0.169
UniGenX(100M) 0.38 65 39 35 0.045 0.194
Table 7: Conditional generation results for five quantum properties (units specified), evaluated using the EDM setting from [61]. Numbers represent the mean absolute error (MAE) between conditional and oracle-predicted properties [65]. The QM9 row indicates the baseline MAE of the EGNN model, used to obtain the oracle-predicted properties for generated samples, compared to QM9 property labels.
Table 6 presents the LDM sampling results for conditional generation as introduced in [62], where UniGenX achieves state-of-the-art (SOTA) performance across all five properties. However, we believe there is an issue with the evaluation results provided by [62]. Following [61], when we used another portion of the QM9 dataset to assess the performance of the classifier from [62], we found that the values for $\epsilon_{H}$, $\epsilon_{L}$ and $\Delta\epsilon$ were approximately 27 times larger than those reported in [61]. Given that $1Ha=27.2114eV$, we suspect there is a unit discrepancy in [62]. While the unit in their paper is listed as meV, we believe the correct unit should be mHa. We highlight this error here and have corrected it in our table.
Table 7 presents the conditional generation results using EDM sampling, as introduced in [61]. UniGenX achieves state-of-the-art (SOTA) performance on four out of six properties. Notably, for the property $\mu$, our model demonstrates a significant improvement of $53.6\%$ over the previous SOTA [70]. However, performance on $\alpha$ and $C_{v}$ is slightly lower compared to [69]. This can be attributed to two factors: first, [69] employs separate models for each property, whereas UniGenX uses a unified model, potentially leading to performance trade-offs. Second, [69] incorporates post-generation molecular force field optimization, following [62], which leverages domain-specific expert knowledge. Our approach evaluates the model’s direct output without such optimization. While these differences may impact direct comparison, UniGenX still excels in key performance metrics, highlighting its versatility and robustness.
The properties in Table 7 were calculated for generated samples using an EGNN model, consistent with the EDM settings from [61]. The QM9 row indicates the EGNN model’s error estimation relative to the QM9 property labels. Notably, UniGenX’s generated results exhibit error closely approximating the validation model’s error. For instance, the HOMO gap is identical to QM9 (39 vs. 39), and the LUMO gap is lower (35 vs. 36). This suggests that UniGenX’s performance may exceed the validation model’s capacity for accurate error assessment. A detailed discussion of this discrepancy is reserved for future work.
Figure 6: Examples of groundtruth (left) and generated (right) molecular structures
Furthermore, four examples are displayed in Figure 6, in each example on the left is ground truth and on the right is the generated conformation.
MP-20 Carbon-24 MPTS-52
Methods MR(%$\uparrow$ RMSD $\downarrow$ MR(%$\uparrow$ RMSD $\downarrow$ MR(%$\uparrow$ RMSD $\downarrow$
CDVAE[53] 33.90 0.1045 17.09 0.2969 5.34 0.2106
DiffCSP[38] 51.49 0.063 17.54 0.2759 12.19 0.1786
FlowMM[39] 61.39 0.057 23.47 0.4122 17.54 0.1726
UniGenX(100M) $\mathbf{64.74}$ $\mathbf{0.042}$ $\mathbf{29.46}$ $\mathbf{0.2318}$ $\mathbf{32.97}$ $\mathbf{0.0838}$
Table 8: Results of the unification model on material tasks. MR represents Match Rate. UniGenX’s performance surpasses the state-of-the-art, which proves its unified ability across diverse domains.
Dataset Large-scale QM9
[…]
Mean Median Mean Median Mean Median Mean Median
ConfGF[58] 46.23 44.87 0.5171 0.5133 89.21 95.12 0.2809 0.2837
GeoMol[57] 78.28 81.03 0.3790 0.3861 91.05 95.55 0.2970 0.2993
DMCG[36] 90.86 95.36 0.2305 0.2258 $\mathbf{98.34}$ $\mathbf{100}$ 0.1486 0.1340
UniGenX(100M) $\mathbf{91.43}$ $\mathbf{100.00}$ $\mathbf{0.2186}$ $\mathbf{0.0818}$ 92.53 $\mathbf{100}$ $\mathbf{0.1441}$ $\mathbf{0.0818}$
Table 9: Results of the unification model on the small molecular task of GEOM-QM9 using COV-P, MAT-P, COV, and MAT metrics. COV-P, MAT-P, COV, and MAT evaluate molecule coverage and accuracy, as defined by Equations (10)–(13).
The finetuned results for MP-20, Carbon 24, MPTS-52 and GEOM-QM9 results can be found in Table 8 and Table 9. The results show that a 100M unification model UniGenX(100M) still performed better results than the SOTA methods on most indicators. With larger models and more data, our model can scale to achieve better results due to the scalability of the auto-aggressive and diffusion models. Additionally, due to the inclusion of a word loss head, our model is adaptable for language training, making it flexible for general tasks such as language-based design and generation.
5.3 Intergrating with LLM Results
MP-20
Methods MR(%$\uparrow$ RMSD $\downarrow$
UniGenX-NL(1B) 63.10 0.0397
Table 10: Results of UniGenX with capability of natural langauge.
The results shown in 10 demonstrate that our model retains its linguistic capabilities without compromising accuracy, providing a solid foundation for future work, such as integrating language with a unified model for both linguistic and numerical tasks.
6 Conclusion
This work introduces UniGenX, a novel framework for unified generation of symbolic and numerical scientific data. UniGenX integrates an autoregressive model with a conditional diffusion-based generative head, leveraging the autoregressive model’s sequence flexibility and the diffusion model’s numerical precision. A key innovation is a sequentialization scheme that represents diverse scientific data—formulas, coordinates, energies, forces—as a single token sequence. This enables next-token prediction with the autoregressive component, while the diffusion head addresses numerical precision limitations by operating in continuous space and training efficiently on low-dimensional variables. UniGenX achieves significant performance improvements over state-of-the-art diffusion/flow matching models in material generation (10–120% gains on MP-20, Carbon-24, and MPTS-52), surpasses NatureLM, and establishes new state-of-the-art results on most targets in GEOM-QM9 prediction, de novo design on materials, and conditional generation on QM9. Demonstrating high capacity for unified training across domains and natural language prompt generation, UniGenX avoids explicit inductive biases for equivariance/invariance, relying on data augmentation to learn these properties, thus preserving transformer scalability. Future work will extend UniGenX to other scientific data types—proteins, DNA—and tasks—energy/force prediction, aiming towards a unified foundation model for scientific generation, potentially integrating with general domain models.
Appendix A Diffusion Model and Flow Matching
The training targets for diffusion and flow matching models share the same mathematical structure, as stated by Theorem 1 [71].
Theorem 1.
Let $X$ be an integrable random variable. Then, for each $\sigma$-algebra $\mathcal{V}$ and $Y\in\mathcal{V}$, $Z=\mathbb{E}(X|\mathcal{V})$ solves the least squares problem
$\min_{Z_{\theta}\in\mathcal{V}}\|Z_{\theta}-X\|\,,$
[…]
A.1 Relationship Between Diffusion Models and Flow Matching
Diffusion models and flow matching share the same noising procedure, given by Eq. (4), i.e.,
$x_{t}=\alpha_{t}x+\sigma_{t}z,$ (16)
where $x\sim p_{0}$, $p_{0}$ is the data distribution, $z\sim p_{prior}$, $p_{prior}$ is a prior distribution from which samples can be easily drawn. In diffusion models, the prior random variable is a standard Gaussian distribution, i.e., $z=\epsilon\sim N(0,I)$, thus limiting the diffusion process to approach Gaussian noise from the data variable. However, flow matching extends the prior to any distribution, making it a flow from samples of the given distribution to the generated data samples [72, 73]. Rectified flow and stochastic interpolation can also be obtained through similar approaches [74, 23, 75]. A typical training target is to find the velocity that connects the prior and data distributions, $\frac{dx_{t}}{dt}=v$. Hence, the training target is directly derived from Eq. (4) as
$\frac{dx_{t}}{dt}=\dot{\alpha}_{t}x+\dot{\sigma}_{t}z,$ (17)
with the neural network $v_{\theta}(x_{t})$ approximating the velocity $v$ of the flow by
$v=\mathbb{E}\left[\frac{dx_{t}}{dt}\middle|x_{t}\right]=\mathbb{E}\left[\dot{\alpha}_{t}x+\dot{\sigma}_{t}z\middle|x_{t}\right]=\dot{\alpha}_{t}\mathbb{E}\left[x\middle|x_{t}\right]+\dot{\sigma}_{t}\mathbb{E}\left[z\middle|x_{t}\right].$ (18)
The score target, denoising model target, and $\epsilon$ target for flow matching with the Gaussian noise can also be derived from Eq. (7) and direct calculation, as discussed in Appendix A.2. In this work, we primarily focus on the combination of autoregressive models and diffusion models, as the combination with flow matching is similar to that with diffusion models. Therefore, this investigation is left for future discussion.
A.2 Connections of the Training Targets
When the random variable $z$ is a standard Gaussian random variable, i.e., $z=\epsilon$, we can express the connection between the diffusion model training targets and the training target $v$ in flow matching. From the property of conditional expectation, $x_{t}=\mathbb{E}[x_{t}|x_{t}]$, one can obtain
$x_{t}=\mathbb{E}[x_{t}|x_{t}]=\mathbb{E}[\alpha_{t}x+\sigma_{t}\epsilon|x_{t}]=\alpha_{t}\mathbb{E}[x|x_{t}]+\sigma_{t}\mathbb{E}[\epsilon|x_{t}].$ (19)
From diffusion models, $x_{\theta^{*}}=\mathbb{E}[x|x_{t}]$ gives the training target of the denoising model, and $\epsilon_{\theta^{*}}=\mathbb{E}[\epsilon|x_{t}]$ gives the training target of the $\epsilon$ model. From Eq. (7), we have
$s(x_{t})=-\frac{x_{t}-\alpha_{t}\mathbb{E}[x|x_{t}]}{\sigma_{t}^{2}}.$ (20)
A direct calculation according to Eq. (18), (19), and (20) yields the connections of the diffusion targets $s_{\theta^{*}}$, $\epsilon_{\theta^{*}}=\mathbb{E}[\epsilon|x_{t}]$, and $x_{\theta^{*}}=\mathbb{E}[x|x_{t}]$ with the flow matching target $v$ as
$s_{\theta^{*}}$ $=\frac{\alpha_{t}v-\dot{\alpha}_{t}x_{t}}{\dot{\alpha}_{t}\sigma_{t}-\alpha_{t}\dot{\sigma}_{t}}\sigma_{t}^{-1};$ (21)
$\epsilon_{\theta^{*}}$ $=\frac{\dot{\alpha}_{t}x_{t}-\alpha_{t}v}{\dot{\alpha}_{t}\sigma_{t}-\alpha_{t}\dot{\sigma}_{t}};$
$x_{\theta^{*}}$ $=\frac{\dot{\sigma}_{t}x_{t}-\sigma_{t}v}{\dot{\sigma}_{t}\alpha_{t}-\sigma_{t}\dot{\alpha}_{t}}.$
Appendix B Pseudocode
We provide the overall algorithm pseudocode, which includes the forward and sample processes for the model, as well as the forward and sample processes for diffloss. The algorithms are listed as follows.
Algorithm 1 UniGenX Forward Procedure
1:
2:$w_{gt}$: Ids of input tokens in the input sentences
3:$v_{gt}$: Numeral values in the input sentences
4:$m_{val}$: Masks for values
5:$m_{pad}$: Padding Masks
6:$n$: Multiplication times of DiffLoss
7:$w$: Loss weight
8:Initialize $x_{input}$ in the shape of $w_{gt}$
9:# Embed the inputs
10:$x_{input}[\neg m_{v}]\leftarrow$ Embedding($w$)
11:$x_{input}[m_{v}]\leftarrow$ Linear($v$)
12:# Pass through the AR Model
13:$h_{rep}\leftarrow$ Transformer_Decoder($x_{input},m_{pad}$)
14:$w_{logits}\leftarrow$ Linear($h_{rep}[\neg m_{val}]$)
15:# Compute the diffloss and wordloss
[…]
18:return $l_{d}+w*l_{w}$
Algorithm 2 DiffLoss Forward Procedure
1:
2:$v_{gt}$: Ground truth numeral values
3:$h_{rep}$: Hidden representations for conditional diffusion
4:$M$: Times of Multiplication of DiffLoss
5:Initialize $\leftarrow$ SimpleMLPAdaLN(), $\leftarrow$ GaussianDiffusion()
6:# Apply DiffMul
7:$v\leftarrow v_{gt}.$repeat($M$)
8:$h\leftarrow h_{rep}.$repeat($M$)
9:# Sample Different Timesteps for Diff
[…]
12:# Apply Diffusion
[…]
14:return $l$
Algorithm 3 GaussianDiffusion Training Loss
1:
2:net: Neural network for predicting noise and variance
3:$x_{0}$: Hidden representations for conditional diffusion
4:$t$: Timesteps for Diffusion
5:$c$: Condition for Diffusion
6:Sample $\epsilon$ from $\mathcal{N}(\bf{0},\bf{I})$
7:$x_{t}\leftarrow\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-{\overline{\alpha}_{t}}}\epsilon$
8:# Predict Noise and Variance using
9:$\hat{\epsilon},\hat{\sigma}=\texttt{net}(x_{t},t,c).\texttt{split}()$
10:# Compute L2 loss and VLB loss
11:$L_{mse}\leftarrow\texttt{L2Loss}(\hat{\epsilon},\epsilon)$
[…]
Algorithm 4 UniGenX Sampling Procedure
1:
2:$w_{in}$: Ids of input tokens in the initial sentences
3:$v_{in}$: Numeral values in the initial sentences
4:$m_{val}$: Masks for values
5:$m_{pad}$: Padding Masks
6:$\texttt{Initialize}~{}w\leftarrow w_{in},v\leftarrow v_{in}$
[…]
8:     # Prepare for next word/value prediction
9:     $n\leftarrow\texttt{prepare\_inputs\_for\_generation}(w,v,m_{val},m_{pad})$
10:     # Embed the inputs
11:     $x_{input}[\neg m_{v}]\leftarrow$ Embedding($n_{w}$)
12:     $x_{input}[m_{v}]\leftarrow$ Linear($n_{v}$)
13:     # Pass through the AR Model
14:     $h_{rep}\leftarrow$ Transformer_Decoder($x_{input},m_{pad}$)
15:     $w_{logits}\leftarrow$ Linear($h_{rep}[\neg m_{val}]$)
16:     # Compute the word/value
[…]
18:     $y_{w}\leftarrow\texttt{multinomial}(\texttt{softmax}(scores))$
19:     # Use Diffloss Sampling Procedure in Alg5
[…]
21:     if $n_{nxt}$ is a value then
22:         $v.\texttt{append}(y_{t})$
23:     else
24:         $w.\texttt{append}(y_{w})$
25:     end if
26:end while
27:return $w,v$
Algorithm 5 Diffloss Sampling Procedure
1:
2:$h_{rep}$: Hidden representations for conditional diffusion
3:$T=200$: Number of sampling steps
4:# Initialize noise, model kwargs and diffusion
[…]
6:$model\_kwargs\leftarrow\texttt{dict}(c=h_{rep})$
[…]
8:$r\leftarrow\texttt{diff}.\texttt{p\_sample\_loop}(\texttt{net},y,model\_kwargs)\textbf{ \# Predicted denoised sample}$
9:return $r$
Appendix C More Details about Dataset Processing
C.1 Dictionary for Diverse Domains and Tasks
In UniGenX, we utilize a straightforward vocabulary dictionary for representing SMILES strings and material formulas, without any specialized design. Each atom type from the periodic table and special characters in SMILES, such as ‘C’, ‘Cu’, ‘(’, ‘)’, ‘=’, are treated as individual tokens. This approach enables UniGenX to accurately parse atoms and infer two-dimensional SMILES information through character-wise attention on special characters. Consequently, UniGenX can be readily extended to other scientific domains by simply expanding the dictionary to include additional special characters, such as amino acid names for proteins and nucleic acid names for DNAs and RNAs.
The vocabulary dictionary also incorporates special tokens, which are crucial for UniGenX’s flexibility as a next-token prediction model across diverse domains. These tokens serve as domain indicators, such as $<$bof$>$ for material formulas and $<$bos$>$ for molecule SMILES. Extending UniGenX to other domains, like proteins and DNA, requires adding corresponding tokens, such as $<$boa$>$ and $<$bod$>$. This design empowers UniGenX to handle various domains within a unified sequence representation. Thus, these domain-specific special tokens are integral to the vocabulary dictionary.
Furthermore, special tokens facilitate different tasks. For structure generation, $<$boc$>$ denotes coordinates. For conditional generation of sequences and structures, property tokens like $<$bulk$>$ for bulk modulus and $<$band$>$ for band gap are used. These tokens, followed by property values, precede the molecule sequence and coordinates. This allows UniGenX to generate both sequences and structures conditioned on specific properties. Therefore, the special tokens in the vocabulary dictionary provide UniGenX with flexibility across domains and tasks.
C.2 Materials
The pretraining dataset comprises structures from the Materials Project [76], NOMAD [77], and OQMD [78, 79], widely used open-source databases for materials. We eliminated duplicates in the merged dataset, filtered out data with chemical formulas failing SMACT validation, and removed structures with formulas matching those in the MP-20, Carbon-24, and MPTS-52 test sets used for evaluation. The final pretraining dataset contained approximately 5.5M samples.
We evaluated UniGenX on three common material benchmark datasets: MP-20, Carbon-24, and MPTS-52. MP-20 consists of 45,231 stable inorganic materials from the Materials Project [76], representing most experimentally synthesized materials with up to 20 atoms per unit cell. Carbon-24 [80] includes 10,153 carbon materials, each with 6 to 24 atoms per unit cell. MPTS-52, an extension of MP-20, is a more challenging dataset comprising 40,476 structures with up to 52 atoms per unit cell, ordered by their earliest publication year. The dataset splits followed those described in [53] and [38].
To ensure data consistency, we sorted atoms and coordinates according to predefined rules. Coordinates were represented as fractional coordinates.
MP-20 MPTS-52 Carbon-24
Train 27,136 32,380 6,091
Valid 9,047 2,032
Test 9,046 8,096 2,030
Table S1: Size of Material Dataset
The parsed data format is as follows:
<bos> [n * sites] <coord> [3 * lattice] [n * frac_coords] <eos>
C.3 Molecules Conformation Generation
For the small molecule conformation generation task, we utilized the GEOM single-molecule multi-conformation dataset. Following DMCG, we partitioned the dataset into training, validation, and test sets with an 8:1:1 ratio, distributing conformations for each molecule into their respective sets to ensure no overlap. This resulted in 1,374,737, 165,204, and 174,162 conformation data points for QM9 in the training, validation, and test sets, respectively. For the larger GEOM-Drugs dataset, we selected 2,000,024, 100,104, and 100,106 conformation data points for training, test, and validation.
During data processing, we converted molecules to Canonical SMILES and renumbered atoms to align their indices with the SMILES order. Special SMILES characters were directly tokenized. In UniGenX, we represent SMILES information as a sequence, rather than a graph, leveraging the transformer decoder’s attention mechanism to capture all inherent information. To address potential mismatches between atom order in SMILES and coordinates, we used Canonical SMILES to align these orders. All information is then represented as a linear sequence input to the model, enabling UniGenX to jointly train SMILES and coordinates, improving molecule understanding compared to sequence-to-structure-only training. The parsed data format is as follows:
<bos> [cano_smiles] <coord> [n * coords] <eos>
C.4 Unification task
We use the mixture of material pretraining dataset and QM9 mentioned above to pretrain UniGenX. By adjusting the ratio of the two datasets, the material dataset loops twice when the small molecule dataset loops once.
The parsed data format is as follows:
<bos> <mat> [n * sites] <coord> [3 * lattice] [n * frac_coords] <eos> <bos> <mol> [cano_smiles] <coord> [n * coords] <eos>
C.5 Conditional Generation of Molecules
For the Conditional Molecule Generation task, we adhered to two settings, as described in [62] and [69]. Both settings utilize the QM9 dataset (approximately 137k samples), partitioning the training data into two halves: one for generator training and the other for classifier training. Consistent with [62], we employed a random seed of 42 for data splitting and utilized their classifier for testing, ensuring no overlap. Unlike [69], we did not perform property normalization. Instead, property values were treated as numeric values and transformed into three-dimensional vectors, which were directly appended to the input sequence. Notably, we trained a unified model, using the same architecture for all properties, a departure from previous methods. Furthermore, unlike graph-based models, our approach does not require pre-sampling atom numbers, as the special token <coord> controls SMILES generation. Finally, our model employs a simple vocabulary dictionary, treating all SMILES characters, including special characters, as individual tokens (see Appendix C.1 for details).
The parsed data format is as follows:
<bos> <prop> [prop_val] [cano_smiles] <coord> [n * coords] <eos>
C.6 Instruction Tuning with Natural Language Model
To train our model to generate crystal structures from natural language instructions, we constructed a dataset using the following format:
<bos>Instruction: Generate a structure based on the provided composition , [n * sites].
Response: [3 * lattice] [n * frac_coords]<eos>
The <bos> and <eos> tokens delineated the start and end of the sequence, respectively, while the Instruction: and Response: labels clearly separated the input and output. By training on this structured data, the model learned to associate textual instructions with corresponding crystal structure representations, effectively bridging the gap between natural language and structural data.
Appendix D Computing Resource Usage
Table S2 details the GPU computational resource usage for various tasks, including task type, dataset, model size, GPU specifications, and computational cost (GPU days). All tasks were trained using float32 precision, as reflected in the table. Tasks were trained using either A100 (80G) or MI300X (192G) GPUs, with computational costs ranging from 0.9 GPU days (Carbon-24) to 191 GPU days (Unified Pretraining). While bf16 precision can achieve an approximate 4x training speedup without compromising accuracy, these results are not included in the table.
Task Data Model GPU Cost (GPU Days)
Material Pretrain UniGenX(400M) MI300X (192G) 109
MP-20 UniGenX(400M) A100 (80G) 2.2
MPTS-52 UniGenX(400M) A100 (80G) 2.6
Carbon-24 UniGenX(400M) A100 (80G) 0.9
Molecule Large-QM9 UniGenX(400M) A100 (80G) 26.6
Large-Drugs UniGenX(100M) MI300X (192G) 128
Conditional Mol QM9 UniGenX(100M) A100 (80G) 12
Unified Pretrain UniGenX(100M) MI300X (192G) 191
Table S2: GPU Usage for Different Tasks
Appendix E Ablation Study
To elucidate the contribution of each component within our model, we performed a series of ablation studies. By systematically removing or altering key components, we aimed to evaluate their individual and combined impact on overall model performance. The core components examined were the data parser, backbone, and diffloss, responsible for data preparation, word and condition generation, and structure generation, respectively. In our ablation studies, we selectively removed or modified these components to analyze their contribution to the final performance.
E.1 Data Parser
In this section, we investigate the impact of modeling the space group (sg) on the performance of the model. Sg train refers to whether the space group information is modeled during the training process, with the aim of evaluating its contribution to the model’s prediction accuracy.
We kept the diff mul and res blocks parameters constant (at 16 and 12, respectively) and conducted a comparative experiment on whether to model the space group.
Figure S1: The distribution of the frequency of the space group in training data is imbalanced.
The experimental results shown in S3 indicate that the model performs slightly better without modeling the space group than when the space group is modeled, due to the imbalanced space group distribution as shown in S1. This may suggest that in this particular task, the space group information does not significantly help to improve the match precision and may even introduce some unnecessary complexity, limiting further enhancement of model performance.
Hence, we can conclude from this ablation study that although space group information might be beneficial for model performance in certain tasks, disabling sg allows the model to learn spatial features more freely in the current task, thereby improving overall performance. In subsequent ablation studies, unless specifically mentioned, the space group will not be modeled.
Model Diff mul Res blocks Sg Match rate %
UniGenX(100M) 16 12 + 54.62
UniGenX(100M) 16 12 - 56
Table S3: Comparison of Model Performance with and without sg
E.2 Model Size
In this section, we conducted detailed experiments to investigate the impact of model size on the performance of the backbone. Specifically, we compared the performance of UniGenX(100M) and UniGenX(400M), and the experimental results are presented in S4. Meanwhile, different model size configurations are in S5
In the experiments, the main differences between UniGenX(100M) and UniGenX(400M) lie in the number of hidden layers and the number of attention heads. The experimental results show that as the model size increased from 100M to 400M parameters, there was an improvement in model performance. Specifically, the Match score for UniGenX(100M) was 57, while the Match score for UniGenX(400M) increased to 59.35. This indicates that increasing the number of hidden layers (from 6 to 24) enhanced the model’s feature representation capability and its ability to capture complex structures, under the same attention head configuration.
These experimental findings validate the advantages of larger models in complex tasks, particularly in terms of their ability to better extract and integrate information within multi-layered deep structures and rich contextual representations. However, it is important to note that while performance improvements are achieved, the increase in model size also brings higher computational costs. Therefore, in practical applications, balancing performance against computational expense is an important consideration.
The experiments consistently demonstrate that increasing model parameters, especially the number of hidden layers, has a positive impact on match precision.
Model Diff mul Res blocks Match rate %
UniGenX(100M) 16 12 57
UniGenX(400M) 16 12 59.35
Table S4: Backbone parameters and performance.
Model UniGenX(100M) UniGenX(400M)
hidden size 1024 1024
intermediate size 4096 4096
num hidden layers 6 24
num attention heads 16 16
num key value heads 16 16
Table S5: Comparison of UniGenX(100M) and UniGenX(400M) model configurations
E.3 Diffloss
E.3.1 Diffmul & Resblocks
In the ablation study of the diff module, we experimented with two key parameters in the module, the diffmul and the resblocks, to assess their impact on model performance.
The diffmul, short for diffusion multiplier, represents the number of time points selected along the path where the model pushes the noise distribution forward to the data distribution. It controls the model’s ability to extract features along the diffusion path. We chose two different values, 4 and 16, to observe the impact of smaller and larger feature extraction capabilities on the final results. As shown in S6, it is obvious that the model performs better with a diffmul value of 16 compared to a value of 4. For example, with res blocks fixed at 3, increasing the diffmul from 4 to 16 improved the Match score from 53.31 to 56.06. Similarly, with res blocks at 12, the Match score for a diffmul of 16 (54.62) was higher than that for a diffmul of 4 (48.98). This suggests that a greater feature extraction capability performs better in handling complex structures.
Resblocks refer to the number of residual blocks, which are used to enhance the model’s expressive power. We compared the performance between models with 3 and 12 Resblocks, observing mixed effects on performance. With space group modeling, a model size of 100M, and a Diffmul of 4, increasing the number of Resblocks from 3 to 12 led to a decrease in Match Rate from 53.31% to 48.98%. However, with a Diffmul of 16, the performance difference between 3 and 12 Resblocks was marginal (56.06% vs. 54.62%). On the other hand, when increasing the model size from 100M to 400M and not modeling the space group, increasing Resblocks from 3 to 12 improved performance from 54.78% to 59.35%. These results suggest that while increasing the number of Resblocks helps the model handle more complex input data, the performance gains may plateau in certain conditions.
In summary, the experimental results demonstrate that increasing the values of diffmul and resblocks significantly contributes to enhancing the overall performance of the model.
Model Diffmul Resblocks Sg Match rate %
UniGenX(100M) 4 3 + 53.31
UniGenX(100M) 16 3 + 56.06
UniGenX(100M) 4 12 + 48.98
UniGenX(100M) 16 12 + 54.62
UniGenX(400M) 16 3 - 54.78
UniGenX(400M) 16 12 - 59.35
Table S6: Experimental Results of Diffloss Module Configurations
E.3.2 Scheduler
In this section, we explore the impact of two schedulers on the effectiveness of the model. The two schedulers are the Diffusion scheduler used in this paper and the EDM scheduler, which was first introduced in [45].
The EDM scheduler utilizes a second-order Heun ODE solver for improved sampling accuracy and efficiency. The time step scaling follows a specific schedule based on $(\sigma_{max}^{\frac{1}{\rho}}+\frac{1}{N-i}(\sigma_{\text{min}}^{\frac{1}{\rho}}-\sigma_{\text{max}}^{\frac{1}{\rho}}))^{\frac{1}{\rho}}$, allowing for smooth transitions during the diffusion process. EDM employs a flexible network architecture, enabling compatibility with various models. Key preconditioning techniques include skip scaling, defined as $\sigma\cdot\sigma_{\text{data}}/(\sigma_{\text{data}}^{2}+\sigma^{2})$, along with tailored input and output scaling formulas that optimize noise and signal propagation. The noise distribution is modeled as $\ln(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$. Loss weighting is also balanced with the formula $(\sigma_{\text{data}}^{2}/(\sigma_{\text{data}}^{2}+\sigma^{2}))^{2}$. More details can be found in [45].
EDM parameters
We provide three images showing the Match rate under different std, sig, and mean values in S2. The default model setting is UniGenX(100M) with 12 resblocks and 16 diffmul. Here is a summary of the main trends:
Impact of the mean value:
  • The match rate tends to be higher, typically maintaining above 0.58 to 0.6, when the mean value fluctuates between -1.4 and -0.8. A relatively lower mean value seems to correlate with a higher match rate.
  • When the mean value decreases to -1.6, the match rate sometimes drops slightly but the overall performance remains relatively stable.
  • Larger mean values (e.g., -0.8) usually perform better in terms of match rate, and as the mean value becomes smaller (e.g., -1.6 or lower), the match rate may gradually decrease.
Impact of std (standard deviation):
  • The match rate usually remains stable when std fluctuates between 1 and 1.7. The smaller the std value (e.g., 1), the match rate tends to be slightly higher.
  • When the std variation is small (e.g., from 1 to 1.5), fluctuations in the match rate are not significant, suggesting that std has a minimal impact on the match rate within this range.
  • As std further increases to 1.7, the match rate begins to decline slightly, but not significantly.
Impact of sigma value:
  • The closer the sig value is to 1, the better the match rate typically performs, often fluctuating between 0.58 and 0.6 when the value is 1.
  • When the sig value is 1.3, the match rate sometimes decreases slightly but still maintains a relatively high level.
  • Lower sig values (e.g., 0.2 or 0.5) are usually accompanied by lower match rates, indicating that lower sig values may inhibit the model’s matching ability.
Multifactorial influence:
  • For the combination of mean = -1.4, std = 1.2, and sig = 1, the match rate can reach a peak of 0.6024.
  • As std and sig increase to higher values, the match rate typically declines, e.g., when sig increases to 1.5, there is a slight decrease in the match rate.
  • A combination of particularly low sig and larger mean (e.g., mean = -1.6, sig = 0.2) results in a sharp decline in the match rate.
In summary, the data suggest that moderate mean and std values combined with a higher sig value can enhance the match rate, while more extreme combinations of mean and sig tend to lower the match rate.
Figure S2: Comparison between mean, std and sigma in EDM
DDPM vs. EDM
We compare Diffloss and EDMloss in S7, and the analysis of the table is as follows.
Despite the fact that the EDM scheduler slightly outperforms the Diffusion loss in its best results, the values of key parameters such as sig, mean, and aug for conditional diffusion are challenging to determine accurately. This introduces a large number of parameter combinations in EDM, requiring extensive hyperparameter tuning. Moreover, after incorporating data augmentation, we observed that the performance gap between EDM and DDPM training strategy becomes minimal. In this work, we mainly focus on the combination of the auto-regressive model and the diffusion head in a unified generation across different tasks and different domains. Hence, here we adopt the DDPM training strategy for its simplicity, requiring fewer hyperparameters to be tuned. To achieve better performance, an investigation of combining different setups of other diffusion strategies or flow matching models is also supported but not included in this work.
Model Diffmul Resblocks Scheduler Augmentation MR %
UniGenX(400M) 16 12 DDPM - 59.35
UniGenX(400M) 16 12 EDM - 59.08
UniGenX(400M) 16 12 DDPM + 63.88
UniGenX(400M) 16 12 EDM + 63.00
Table S7: Ablation study on Scheduler and Augmentation.
E.4 Data Augmentation
In this section, we investigate the impact of data augmentation on model performance. The augmentation techniques employed include unit cell translation and rotation. For unit cell translation, a random three-dimensional vector is added to the fractional coordinates of each atom, and the result is taken modulo 1 to maintain periodicity. Conversely, unit cell rotation involves randomly rotating the lattice coordinate system while keeping the atomic coordinates fixed. Algorithm 6 presents the pseudocode for material data augmentation.
Algorithm 6 Rotation & Translation Augmentation
1:
2:$l$: Lattice
3:$x$: Fractional Coordinates
4:# Apply Rotation
[…]
6:$l\leftarrow R\cdot l$
7:# Apply Translation
8:$t\sim\mathcal{N}(\vec{0},\bf{I}_{3})$
9:$x\leftarrow(x+t)\mod 1$
10:return $l,x$
As shown in Table S7, incorporating augmentation significantly enhances performance for both schedulers, resulting in an approximate 4-point increase in match score.
E.5 Pretraining or Training from Scratch
In this section, we present an ablation study comparing the impact of pretraining versus training from scratch across three datasets: MP-20, Carbon-24, and MPTS-52. Table S8 summarizes the results, detailing Match Rate (MR) and Root Mean Square Deviation (RMSD).
The pretrained base model exhibits moderate performance across the datasets, achieving a Match Rate (MR) of 60.48% on MP-20, 3.15% on Carbon-24, and 33.61% on MPTS-52, with corresponding RMSD values of 0.0568, 0.1928, and 0.1086, respectively.
Training the model from scratch yields a slight improvement in MR for MP-20, reaching 63.88%, but demonstrates mixed results on Carbon-24 and MPTS-52, with MR decreasing to 27.09% and 29.09%, respectively.
Conversely, finetuning the pretrained model on each dataset results in superior performance, with substantial enhancements in MR across all three datasets. On MP-20, the finetuned model achieves an MR of 67.01%. Similarly, on Carbon-24 and MPTS-52, the finetuned model achieves MR values of 30.05% and 38.65%.
These findings suggest that pretraining followed by dataset-specific finetuning is a more effective strategy than training from scratch, particularly for complex datasets like MP-20 and MPTS-52. Notably, even when trained from scratch, UniGenX outperforms FlowMM on these three benchmarks, demonstrating the inherent advantages of UniGenX.
MP-20 Carbon-24 MPTS-52
Methods MR(%$\uparrow$ RMSD $\downarrow$ MR(%$\uparrow$ RMSD $\downarrow$ MR(%$\uparrow$ RMSD $\downarrow$
Base Model 60.48 0.0568 3.15 $\mathbf{0.1928}$ 33.61 0.1086
Train from Scratch 63.88 0.0598 27.09 0.2264 29.09 0.1256
Finetune on Each $\mathbf{67.01}$ $\mathbf{0.037}$ $\mathbf{30.05}$ 0.2286 $\mathbf{38.65}$ $\mathbf{0.0657}$
Table S8: Pretraining or Training from Scratch.